Classify the following solvents as either protic o
Answer #1
Concepts and reason

The given solvents are classified into protic and aprotic based on their connectivity of the atoms.

The structure of all the solvents are drawn to check whether the hydrogen atom present in the solvent is bonded to an electronegative atom or not.

Fundamentals

Electronegativity of an atom is defined as the property of an atom to attract the bonded electron pair towards itself.

Protic solvent is the solvent in which the hydrogen atom is attached to an electronegative element. The electronegative element can nitrogen, oxygen or fluorine.

These solvents contain active H+{{\rm{H}}^ + } in their solutions. These solvents participate in hydrogen bonding in the reactions. They serve as acids.

Aprotic solvent is the solvent in which the hydrogen atom is not attached to an electronegative element.

These solvents do not contain active H+{{\rm{H}}^ + } in their solutions. These solvents do not participate in hydrogen bonding in the reactions.

The protic solvents are the ones which have hydrogen atom attached to the electronegative atom.

The protic solvents are as follows:

CH3OHCH3CH2OHCH3COOH(aceticacid)H2O\begin{array}{l}\\{\rm{C}}{{\rm{H}}_{\rm{3}}}{\rm{OH}}\\\\{\rm{C}}{{\rm{H}}_{\rm{3}}}{\rm{C}}{{\rm{H}}_{\rm{2}}}{\rm{OH}}\\\\{\rm{C}}{{\rm{H}}_{\rm{3}}}{\rm{COOH}}\left( {{\rm{acetic}}\,{\rm{acid}}} \right)\\\\{{\rm{H}}_{\rm{2}}}{\rm{O}}\\\end{array}

The aprotic solvents are the ones which have hydrogen atom attached to the electronegative atom.

The aprotic solvents are as follows:

DMFDMSOacetoneacetonitrile\begin{array}{l}\\{\rm{DMF}}\\\\{\rm{DMSO}}\\\\{\rm{acetone}}\\\\{\rm{acetonitrile}}\\\end{array}

Ans:

The given solvents are classified into protic and aprotic solvents as follows:

Protic
Aprotic
|СН,ОН
DMF
DMSO
acetonitrile
сн;CH,он
acetic acid
H20
acetone